HUBBLE SPACE TELESCOPE DAILY REPORT #2190 PERIOD COVERED: 0000Z (UTC) 08/18/98 - 0000Z (UTC) 08/19/98 Daily Status Report as of 231/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD/WF/PC-2/NIC/1/2/3 7174 (LBDS 53W091: A Cluster of Old, Red Galaxies) The Space Telescope Imaging Spectrograph (CCD), the WF/PC-2 and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the cluster of old red galaxies known as LBDS53W091. This distant cluster { at Z=1.55 } is representative of the oldest stellar systems we can glimpse at their "middle age"; and we are interested in their morphologies. A wavelength calibration image was also taken. The STIS portion of this nine-hour proposal was affected by the STIS being in suspend from the beginning of the proposal until it was recovered at 230/1104Z. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.3 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell V4169-SGR in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.4 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of NIC/2 7835 (A Search for Superplanets Embedded in Beta Pic & Vega-like Circumstellar Disks) The Near Infrared Camera (Camera Two) was used to make observations of the circumstellar disk around star HD27290 in order to search for superplanets. The discovery of a circumstellar disk orbiting Beta Pictoris as well as the discovery of other similar IR-excess Vega-like systems have provided compelling evidence that our solar system, or at least many of the processes which are thought to have created it, may be far from unique. The STIS portion of this two-hour proposal was affected by the STIS commanded to suspend as described in 2.3. Otherwise, The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed S/C/STIS/MA2 7647 (STIS NUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spectrograph (MA2) was to be used to make near-ultraviolet (NUV) flat calibration images. This program was to obtain NUV-MAMA flat-field observations with the D2 lamp for the construction of on-orbit D-flats for G230M. A dark calibration image was also to be taken. This proposal was affected by the STIS being commanded to suspend (see 2.3). 1.7 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was to be used to make MAMA dark measurements. These darks were to be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. This proposal was affected by the STIS being commanded to suspend (see 2.3). 1.8 Completed STIS/CCD/MA1/MA2 7656 (Spectroscopic and Imaging Sensitivity) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was to be used to make sensitivity measurements for all supported CCD imaging and spectroscopic modes. It is run once a year. This proposal was affected by the STIS being commanded to suspend (see 2.3). 1.9 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.11 Completed Two Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was to be used to make high galactic latitude observations take in parallel with the prime science observations. This proposal was affected by the STIS being commanded to suspend (see 2.3). 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 23 Successful : 23 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. Per ROP NS-5, the SI C&DH errors were reset three times. Operations supported the IP spacecraft test with TDRS East from 230/1545Z until 230/1618Z. A total of 838 commands were transmitted in the COUNTCHECK mode. No problems were experienced except duting the transmission of a second FIVEBLCK procedure, at 230/1611Z, SCD received an invalid sync code. All subsequent commanding executed nominally. The error was discussed with the IP developers. Further testing will be conducted today. The STIS instrument was recovered from suspend at 230/1104Z per the direction of an operations request. Before the recovery, the proposal mentioned in 1.2 was affected. The proposal described in 1.4 then was performed nominally. The erroneous commanding, an invalid MAMA optical mode selection, which caused the suspend yesterday was due to occur again at approximately 230/1800Z, so operations personnel commanded the STIS to suspend at 230/1730Z. The STIS flight software error counter was cleared at 230/1733Z per ROP NS-12, and the command error counter was cleared at 230/1727Z per ROP SR-1. These actions affected the proposals described in 1.5, 1.6, 1.7, 1.8, and 1.12 as well as two iterations of STIS proposal 7948 which occur during the period of the next report. STIS was again recovered from suspend at 231/0151Z. It is currently operating nominally. The erroneous commanding is not scheduled to occur during the rest of this operational week. During the event described in the preceding paragraph, at 230/2016Z, real-time operations returned to the MOR from the SAMS back-up facility. Upon the return, operations was unable to transmit data to PODPS. The system manager was informed and is expected to correct the problem this morning. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH